Each of the four independent situations below describes a finance lease in which annual lease payments are payable at the beginning of each year. The lessee is aware of the lessor's implicit rate of return.

Situations 1 2 3 4
Lease term (years) 3 4 5 6
Lessor's rate of return 9% 8% 7% 6%
Fair value of lease asset $60,000 $90,000 $92,000 $85,000
Lessor's cost of leased asset $55,000 $75,000 $78,000 $85,000
Residual Value:
Estimated fair value 0 $20,000 $16,000 $14,000
Guaranteed fair value 0 0 $16,000 $20.000

Calculate the amount the lessee would record as a right-of-use asset, for each of the above situations. Round to the nearest dollar. Format ($XX,XXX)

Situation 1:

Situation 2:

Situation3:

Situation4:
Answer #1
Lease payments Residual value guarantee PV of lease payments PV of residual value Guarantee Right-of-use Asset/lease liability
Situation 1 23,703 0 60,000 0 60,000
Situation 2 22,735 0 75,300 0 75,300
Situation 3 19,656 0 80,592 0 80,592
Situation 4 14,966 5,000 73,592 4,229 77,821

Situation 1

Fair Value of the asset 60,000
Present value of annuity due factor for n = 3 and i=9% 2.53129
Lease Payments at the beginning of each of the next three years(60,000/2.53129) 23,703

Situation 2

PV of residual value (20,000 * 0.73502) 14,700
Present value factor for n = 4 and i = 8% 0.73502
Present value of annuity due factor for n = 4 and i=8% 3.31212
Lease payments = (fair value – present value of residual value) / Present value of annuity due factor = (90,000-14,700) / 3.31212 22,735
Present value of lease payments = (fair value – present value of residual value) =90,000 - 14,700 75,300

Situation 3

PV of residual value (16,000 * 0.71298) 11,408
Present value factor for n = 5 and i = 7% 0.71298
Present value of annuity due factor for n = 5 and i=7% 4.10019
Lease payments = (fair value – present value of residual value) / Present value of annuity due factor = (92,000-11,408) / 4.10019 19,656
Present value of lease payments = (fair value – present value of residual value) =92,000 - 11,408 80,592

Situation 4

PV of guaranteed residual value (20,000 * 0.70486) 14,097
Present value factor for n = 6 and i = 6% 0.70486
Present value of annuity due factor for n = 6 and i=6% 4.91732
Lease payments = (fair value – present value of guaranteed residual value) / Present value of annuity due factor = (85,000-14,097) / 4.91732 14,966
Present value of lease payments = (fair value – present value of guaranteed residual value) =85,000 - 11,408 73,592
PV of residual value Guarantee =6,000*0.70486 4,229
Right-of-use Asset/lease liability (73,592 + 4,229) 77,821